# Fast offline decoding with local message-passing automata

### v1

@@ Appendix C Constant slowdown under desynchronization
We begin with the bound on $\mathsf{t}_{\sf sim,max}(t)$, which is simpler. For a fixed site $\mathbf{r}$ and time $t$, consider the event $\mathsf{t}_{\sf sim}(\mathbf{r},t)=m$. The number of attempted updates at site $\mathbf{r}$ by time $t$ must be at least $m$. Let $\delta_{k}$ denote the time interval between the $(k-1)$th and $k$th updates at site $\mathbf{r}$, with $k=1,\dots,m$. The event $\mathsf{t}_{\sf sim}(\mathbf{r},t)=m$ implies the event $S_{m}=\sum_{k=1}^{m}\delta_{k}\leq t$. Fixing a constant $\lambda>2e$ and letting $m_{t}$ denote the smallest integer larger than $\lambda t/\mu$,
@@ Appendix D Comments on PDE-based approaches
Let us briefly recall the general construction of [13, 14]. We write the classical variables on each site as $\mathcal{H}_{\sf cl}={\sf span}\{|\phi_{\mathbf{r}},\sigma_{\mathbf{r}}\rangle\}$, where $\phi_{\mathbf{r}}$ is a rotor field valued in $[0,2\pi)$, [^45] Although a truncation to a field which only stores $\log\log L$ different values is permissible for the same reasons as in the main text. and take $\phi$ to be updated according to some appropriate discretization of a differential equation in which anyons act as sources. Letting $\phi_{\mathbf{r},t}$ be the value of $\phi_{\mathbf{r}}$ at time step $t$, the automaton rule $\mathcal{A}$ is then designed to ensure that
@@
From the perspective of this work, a more serious problem with choosing $\mathcal{L}=\partial_{t}-f(\nabla)$ (for any function $f$) is the self-interactions of anyons with the fields they themselves produce. When an anyon moves, it leads to a violation of $\mathcal{L}\phi=\mathsf{A}$ when it “leaves behind” the divergence of $\phi$ that it sourced. If $\phi$ updates at an $O(1)$ speed $v$—meaning that $v$ updates of $\phi$ occur for each application of the feedback $\mathcal{F}$—this anyon will then be attracted to the lattice site it most recently occupied, preventing anyon matching from occurring. The detailed analysis of Ref. [13] showed that this problem can be overcome only by taking $v$ to diverge at least polylogarithmically with $L$, so that $\phi$ very rapidly re-equilibrates after each anyon movement.
@@ E.1 One dimension
Fix a number $10<q<$ 1 and in integer $m>1$. The probabilistic automaton rule $\mathcal{A}_{m,q}^{(1)}$ uses a collection of i.i.d random variables $\xi_{r}$, each drawn from the uniform distribution on $[0,1]$ and refreshed at each time step, and is defined as
@@
Consider a single well-isolated anyon fixed at the origin $r=0$, and examine the expected particle density $\rho(r>0)=\langle n^{+}_{r}\rangle$ at long times $t\gg r$. The annihilation in Def. E.1 leads to an equilibrium density that decays with $r$ in a manner fixed by $m$. For all choices of $m>0$, numerics show that $\rho(r)\sim r^{-\alpha_{m}}$, where the exponent $\alpha_{m}$ decreases with $m$; see Fig. 19 for several choices of $m$ [^48] This can be anticipated from mean field: there we would (schematically) have $\partial_{t}\rho\sim\partial_{r}(v+D\partial_{r})\rho-\rho^{m}$, with a power law scaling suggesting $\rho\sim r^{-\alpha_{m}}$ with $\alpha_{m}=1/(m-1)$; the real power law is (as expected) slower than this. For example, at $m=2$ one finds $\rho(r)\propto\frac{1}{\sqrt{r}}$. This can in fact be shown rigorously using methods developed for analyzing systems undergoing 1D ballistic annihilation [56], but since it is rather tangential to the main point of this paper we will not go through the details.
@@
While we did not explicitly discuss models with power-law interactions in Sec. 5, one may use the same logic to show that decoders with retarded attractive interactions of $1/r^{\alpha>D}$ have thresholds, provided the interactions are set up using the multi-flavor framework of Sec. 3.3. If we use the same feedback operators as in Def. 3.1—so that an anyon at $r$ moves only when $n^{\pm}_{r}\neq 0$—a proof along the same lines does not directly go through, as in this case the erosion of a $(W,\infty)$ cluster takes time $t_{W}=W^{\beta}$ with $\beta>1$. This can be heuristically seen as follows. Suppose a message source is fixed at the origin, and consider a well-isolated anyon at position $W>0$. The anyon undergoes biased motion with a velocity $-v(r)$, where $v(r)\sim\langle n^{+}_{r}\rangle$. If the latter decays as $1/r^{\alpha}$, then
@@ E.2 Higher dimensions
Fix an integer $m>0$ and a number $10<q<$ 1. The message fields $n^{\mathsf{d},c}_{\mathbf{r}}$ are emitted from anyons, and are designed to propagate along ${\mathbf{\hat{c}}}$ in the $\mathsf{d}$ quadrant of the plane. For $\mathsf{d}=1$, the update rules are
@@
$\mathcal{A}_{m,q}^{(2)}(\mathbf{m},{\bm{\sigma}})[n^{1,y}_{(x,y)}]$ =$=\begin{dcases}\mathcal{A}^{(1)}_{m,q}(\mathbf{m}|_{L_{x}},{\bm{\sigma}}|_{L_{x}})[n^{+}_{y}]&n^{1,x}_{(x,y-1)}=0\\ 1&\text{else}\end{dcases}.$
where $L_{x}=\{\mathbf{r}\,:\,r^{1}=x\}$ and similarly for $L_{y}$. In words: $n^{1,c}$ messages are sourced at anyon locations and move along ${\mathbf{\hat{c}}}$ according to the exclusion process of Def. E.1, and $x$-traveling walkers source $y$-traveling walkers. $\mathcal{A}_{m,q}^{(2)}$ updates the fields in the remaining three quadrants analogously, and the full automaton rule is obtained by taking $v$ repetitions of $\mathcal{A}_{m,1}^{(2)}$.
@@
$\rho(\mathbf{r})=\sum_{a,b,c}\langle n^{ab,c}_{\mathbf{r}}\rangle\sim\frac{1}{||\mathbf{r}||_{1}^{\alpha_{m}}}.$ (209)

### v4

@@ Appendix C Constant slowdown under desynchronization
We begin with the bound on $\mathsf{t}_{\sf sim,max}(t)$, which is simpler. For a fixed site $\mathbf{r}$ and time $t$, consider the event $\mathsf{t}_{\sf sim}(\mathbf{r},t)=m$. The number of attempted updates at site $\mathbf{r}$ by time $t$ must be at least $m$. Let $\delta_{k}$ denote the time interval between the (k1)(k-1)th and $k$th updates at site $\mathbf{r}$, with $k=1,\dots,m$. The event $\mathsf{t}_{\sf sim}(\mathbf{r},t)=m$ implies the event $S_{m}=\sum_{k=1}^{m}\delta_{k}\leq t$. Fixing a constant $\lambda>2e$ and letting $m_{t}$ denote the smallest integer larger than $\lambda t/\mu$,
@@ Appendix D Comments on PDE-based approaches
Let us briefly recall the general construction of [13, 14]. We write the classical variables on each site as $\mathcal{H}_{\sf cl}={\sf span}\{|\phi_{\mathbf{r}},\sigma_{\mathbf{r}}\rangle\}$, where $\phi_{\mathbf{r}}$ is a rotor field valued in $0,2\pi)$, [^46] Although a truncation to a field which only stores $\log\log L$ different values is permissible for the same reasons as in the main text. and take $\phi$ to be updated according to some appropriate discretization of a differential equation in which anyons act as sources. Letting $\phi_{\mathbf{r},t}$ be the value of $\phi_{\mathbf{r}}$ at time step $t$, the automaton rule $\mathcal{A}$ is then designed to ensure that
@@
From the perspective of this work, a more serious problem with choosing $\mathcal{L}=\partial_{t}-f(\nabla)$ (for any function $f$) is the self-interactions of anyons with the fields they themselves produce. When an anyon moves, it leads to a violation of $\mathcal{L}\phi=\mathsf{A}$ when it “leaves behind” the divergence of $\phi$ that it sourced. If $\phi$ updates at an $O(1)$ speed vv—meaning that $v$ updates of $\phi$ occur for each application of the feedback $\mathcal{F}$—this anyon will then be attracted to the lattice site it most recently occupied, preventing anyon matching from occurring. The detailed analysis of Ref. [13] showed that this problem can be overcome only by taking $v$ to diverge at least polylogarithmically with $L$, so that $\phi$ very rapidly re-equilibrates after each anyon movement.
@@ E.1 One dimension
Fix a number $0<q<1$ and in integer $m>1$. The probabilistic automaton rule $\mathcal{A}_{m,q}^{(1)}$ uses a collection of i.i.d random variables $\xi_{r}$, each drawn from the uniform distribution on $[0,1]$ and refreshed at each time step, and is defined as
@@
Consider a single well-isolated anyon fixed at the origin $r=0$, and examine the expected particle density $\rho(r>0)=\langle n^{+}_{r}\rangle$ at long times $t\gg r$. The annihilation in Def. E.1 leads to an equilibrium density that decays with $r$ in a manner fixed by $m$. For all choices of $m>0$, numerics show that $\rho(r)\sim r^{-\alpha_{m}}$, where the exponent $\alpha_{m}$ decreases with $m$; see fig. 20 for several choices of mm. [^49] This can be anticipated from mean field: there we would (schematically) have $\partial_{t}\rho\sim\partial_{r}(v+D\partial_{r})\rho-\rho^{m}$, with a power law scaling suggesting $\rho\sim r^{-\alpha_{m}}$ with $\alpha_{m}=1/(m-1)$; the real power law is (as expected) slower than this. For example, at $m=2$ one finds $\rho(r)\propto\frac{1}{\sqrt{r}}$. This can in fact be shown rigorously using methods developed for analyzing systems undergoing 1D ballistic annihilation [64], but since it is rather tangential to the main point of this paper we will not go through the details.
@@
While we did not explicitly discuss models with power-law interactions in sec. 5, one may use the same logic to show that decoders with retarded attractive interactions of $1/r^{\alpha>D}$ have thresholds, provided the interactions are set up using the multi-flavor framework of sec. 3.3. If we use the same feedback operators as in Def. 3.1—so that an anyon at $r$ moves only when $n^{\pm}_{r}\neq 0$—a proof along the same lines does not directly go through, as in this case the erosion of a $(W,\infty)$ cluster takes time $t_{W}=W^{\beta}$ with $\beta>1$. This can be heuristically seen as follows. Suppose a message source is fixed at the origin, and consider a well-isolated anyon at position $W>0$. The anyon undergoes biased motion with a velocity v(r)-v(r), where $v(r)\sim\langle n^{+}_{r}\rangle$. If the latter decays as $1/r^{\alpha}$, then
@@ E.2 Higher dimensions
Fix an integer $m>0$ and a number $0<q<1$. The message fields $n^{\mathsf{d},c}_{\mathbf{r}}$ are emitted from anyons, and are designed to propagate along $^{\mathbf{\hat{c}}}$ in the $\mathsf{d}$ quadrant of the plane. For $\mathsf{d}=1$, the update rules are
@@
[…]
where $L_{x}=\{\mathbf{r}\,:\,r^{1}=x\}$ and similarly for $L_{y}$. In words: $n^{1,c}$ messages are sourced at anyon locations and move along $^{\mathbf{\hat{c}}}$ according to the exclusion process of Def. E.1, and $x$-traveling walkers source $y$-traveling walkers. $\mathcal{A}_{m,q}^{(2)}$ updates the fields in the remaining three quadrants analogously, and the full automaton rule is obtained by taking $v$ repetitions of $\mathcal{A}_{m,1}^{(2)}$.
@@
[…]